OpenStudy (anonymous):

Medal Is -1 a solution of y^2 -1 = 4y + 3 ?

OpenStudy (campbell_st):

no its not if you substitute it you get \[(-1)^2 - 1 = 4 \times (-1) + 3\]

OpenStudy (campbell_st):

to get a solution you need to rewrite the equation as \[y^2 - 4y - 4 = 0\] then to get a solution you will need the general quadratic formula

OpenStudy (anonymous):

ohh i understand thank you so much